# LLMs and the Human Condition

### v5

@@ Abstract
This paper proposes a model of what LLMs are doing when they impress us with their language skills. The model integrates three established theories of human decision making from philosophy, sociology, and computer science. The paper starts with the collective understanding of reasoning from the early days of AI research - primarily because that model is how we humans think we think. It then describes what is commonly thought of as "reactive systems" which is the position taken by many philosophers and indeed many contemporary AI researchers. The third component to the proposed model is from sociology and extends the reactive approach to communities as “agents” and devolves human intelligence to primarily an ecology of practices. Although not flattering to our modern ego, the approach provides an explanation of apparent “mind reading” in human communication, and explains why a “Generative” Pre-trained Transformer (ChatGPT) might be doing something useful.
@@
For the last few years work in AI has become synonymous with machine learning and in particular neural net architectures for Large Language Models or LLMs. These systems do impressive things, but how they do them is poorly understood. Existing explanations of how they work talk of abstractions and “distributed representations” but such concepts map the known properties of the software into a naive folk psychological understanding of how reasoning works. For 35 years computer scientists have been coming round to the idea that intelligence is somehow grounded in the embodied nature of being Brooks (1991); Wooldridge (2023), and in philosophy can be traced back to Wittgenstein finishing work on The Tractatus Wittgenstein (1921) and his Blue and Brown Books Wittgenstein (1958). The problem has been that, with the growth of computers, many researchers have come from outside the discipline and have not studied the history. The idea that we do things by thinking about them is very appealing and those that do not study the history are bound to repeat the mistake.
The paper starts with a recap of good old fashioned AI (GOFAI). It then describes what Brooks and many philosophers have said about the nature of situated action. It then introduces a third mechanism that was touched on in the 1970s in AI research but which is an idea that is hundreds of years old in what is now sociology. Finally the model being developed is applied to a major issue in Natural Language Understanding (NLU) namely the phenomenon of apparent “mind reading”.
@@ 1 The problem: people read minds
Working on NLU, one sooner or later comes across a phenomenon in which humans appear to do “mind reading”. In what follows the focus is one example but the phenomenon itself is extremely common. The solution is often assumed to be a better (and more detailed) representation of the context and from that perspective this paper proposes representing context, not as things, but as an ecology of practices.
@@ 16 Mother: You want some cereal?
All English speakers by definition Tomasello (2008) will follow what is happening in this conversation and feel there is nothing extraordinary about it. The problem for a reference model of language understanding is that the mother says nothing related to the child’s utterances until line 11. There is no overlap of semantic content - none at all. Of course, as English speakers, we can all say what both the mother and the child want.
This has historically lead researchers interested in conversational agents off into the realm of recognition of intent and modelling theory-of-mind (ToM) Allen et al. (1995); Levinson (2006). Classically this would be modelled in terms of the beliefs, desire and intentions of an agent, the theory being that a rational agent will do what it believes is in its interests Cohen and Levesque (1990), all in terms of good old fashioned symbolic representations.
@@ 2 Artificial Intelligence - a history
The reference model of meaning is generally how people see language as working, and the early AI researchers saw the problem as one of mapping messy natural languages into something more tidy Mel’cuk (1981) – more pure – as God would have made it Eco (1995). For centuries there has been the idea that meaning bottoms out at some kind of semantic primitives Ishiguro (1990) and in this tradition Roger Schank proposed that there was a workable set of primitive verbs or ‘acts’ from which the meaning of a sentence could be based Schank (1972). Examples are PTRANS for “physical transfer” and MTRANS for mental or knowledge transfer, and that machines could translate natural descriptions into (representations of) meaning and figure out consequences from there. The primitives approach to AI didn’t really work but there was also the idea that a formal system of symbol manipulations might map isomorphically onto events in the world. In “Godel, Escher, Bach”, Hofstader Hofstader (1979) gives a beautiful example with his pq- system which is a set of syntactic transforms on strings of ’p’, ’q’, and ’-’. The system produces an infinite set of strings including “-p--q---” and “--p--q----” but not “-p---q--”. The number of ’-’s in produced strings just happens to map onto the number of apples in a box as you add more, or take them out. The symbol ’p’ seems to mean “plus” and ’q’, “equals” because they appear in the system of productions in an isomorphic relationship with our system of adding and subtracting. They have no meaning in themselves without the system in which they participate. Rather than meaning being built up by combining primitive features with innate meaning, the “primitives” get meaning from the structure in which they occur.
@@
Computers are good at this kind of syntactic symbol manipulation, and between that and Schank’s observation that it is hard to think of a thousand things you know a hundred things about, this kind of reasoning formed the basis of thinking about computers and meaning. Today a spreadsheet is a great way to automate reasoning about some things but it turns out there is a wide range of tasks that fail when handled this way. Counting apples in a box might seem easy, but a box might have 15 apples in it according to Waitrose, 25 according to Aldi, and a horse would eat the lot no matter how many are rotten. Modelling the world with referential symbols is hard because the referent can change. If the idea is to map a room before you vacuum it, cats and people tend move and re-set the model. In the early days of robot football the machines would wait for the ball to stop before deciding what to do. Air traffic control has a similar problem. At the time of writing aircraft have been grounded while air traffic controllers sort out why an aircraft, identified by flight number, occurred in two locations. There is something wrong with the system’s representation of the state of the world. The feeling in the late 1980s was that no matter how powerful computers might become in the future, simply getting the data for symbolic representations was going to be problematic. There were projects set up to encode all the data in the world, the CYC project Lenat et al. (1986) being the most explicit, but perhaps there was a better way.
@@
How this pans out for machines is still not clear and Brooks’ robots are often dismissed as being just “insect level intelligence” with robot developers continuing to add symbolic representations in order to support higher level intelligence. Philosophers on the other hand continue to push the limits of what higher level behaviours can be explained in terms of reactive systems Hendriks-Jansen (1996); Ward et al. (2017); Gallagher (2020); Bar-On (2021). The notion of a “reactive system” describes a mechanism used for AI problems, but it is not an interesting “algorithms and data-structures” solution. The ELIZA mechanism takes a set of if-then rules to map input to output, and applied to the DOCTOR script was found to be quite engaging. That same mechanism is used today for a myriad of chatbots but at the time and for many years after, AI researchers would consider it just a trick.
@@ 3 Natural Intelligence - a theory
We humans are not entirely reactive agents because we can, on occasion, think about things like whether I can afford that whiskey using an expression such as “£38 - £44 = …” but we also have another trick. We humans modify the environment to fit with us. In the cooking gnocchi for lunch scenario, none of the things I interact with are naturally occurring – someone made them. What is more, these things were put where I would find them. I might use insect level intelligence to eat lunch, but surely others must have had thoughts (with symbols) in order to create our benign environment.
Recently we have been talking with sociologists who look at human decision making at the macro level. The strong reductionist position is alive and well in sociology and many take the line that collective action is simply made up from the behaviour of individuals making their own decisions - basically that sociology is just psychology on a massive scale in the same way as drunk physicists will claim chemistry is just physics. Social simulation based on the behaviour of individuals is of course entirely possible with today’s computers. There are however problems with this in that collective action does not fit with rational self interest of individuals. Why did Britain leave the EU? And why does Germany want to go to war with a nation that 2 years ago was a key trading partner? At the other end of the “structure and agency” debate wikipedia in sociology is the idea that society consists essentially of structures such as institutions, traditions, and norms, and that individuals are merely acting out their role in those structures. Post The Enlightenment we collectively like to think we can reason about things and make rational decisions that are better than the status quo. And it is of course not very flattering to be told you are merely a cog in a machine, but a key finding of AI research over the last 70 years is of course that rationality is not all it is cracked up to be.
@@
And like an ecology in the natural environment, there is a certain amount of autopoiesis resulting in an ecology of practices being robust. In the animal kingdom, if one species is suddenly removed a gap appears in the resource/consumer space. The gap is either quickly filled by the reminants of that spiecies, something else comes along to fill the space – rabbits replace bilbys – or the resources are appropriated by neighbours in that space. Reduce the number of cats in Sydney, and brown snakes move in to eat the rats and mice ABC (Australia). If the local supermarket closes, I do not starve, but engage another no doubt older hunter/gatherer practice. I wander about attentively until I find indicators of something to eat - an Aldi sign perhaps - which triggers the more modern practice of shopping.
@@ 4 A ’pragmatics first’ model of NLU
We animals behave in accordance with both the reactive framework in which the environment “causes” action, and a framework of practices which evolved through evolutionary pressures on our society. So how does this relate to the mechanism of language as interaction? Language can of course be used to describe practices in terms of things as described above, but using language is also part of practices. In detail, returning to the problem of mind reading, the radical enacivists have claimed that in many cases what looks like reasoning about other minds and ToM is actually the “direct perception” of another’s intent. (see discussion of Hutto in Gallagher (2020)) The aim here is to describe a mechanism for doing that.
The basis of the mechanism is that humans go through life with a set of practices that they know how to do. Rather than sensing the world, forming a representation of the salient parts, and then doing (symbolic) reasoning to produce action that happens to string together to form practices, we imitate the practices of others at a shallow level (see for example Gergely and Csibra Gergely and Csibra (2006)). These practices have evolved at the cultural level. There are certainly species specific behaviours that enable it, but an "ecology of practices" as used here is a product of a culture, that is used to propagate the culture.
How does a head learn practices? The proposal is that a head does good old fashioned unsupervised learning on the raw data of what actions lead to which opportunities. The world, in combination with the human body’s ability to act on it, provides interaction that is consistent enough for regularities to be detected. As a consequence, the mind has a catalog of "sequences of interaction" - of practices - which are coherent enough to provide something with a beginning, a middle, and an end. When the agent/world interaction triggers a practice beginning, the agent can predict that, should the middle follow, the end will occur. If the agent can provide that middle, and somehow wants the outcome, then the agent can do what looks like rational action without using symbolic representations of anything. The child in our example dialog above can want to eat something and, from past experience know that saying "I’m hungry" often results in mother feeding him. We might think that the child means the words he says, and he might, but the act of saying it causes the world to change such that there is no longer a sensation of hunger.
@@
Key to this mechanism is that the practices are largely shared by the community of speakers, and that the practices are largely visible to others. If someone walks up to an ATM machine with a card in their hand, it is fairly likely that they will not start ordering a coffee. This is obvious to us humans – we ‘directly perceive,’ not just the actions, but also the trajectory of which the actions are the start. And it is not just humans. A sheep dog does not need to model the mind of a sheep – the sheep’s observable actions are simply a recognisable part of a practice that sheep perform (Figure 2).
Figure 2: Zak does’t read the Ewe’s mind, but instead sees a nasty head but unfolding.
@@
5 What LLMs are actually doing
The proposal is that the things we write down are to a very large extent descriptions of practices. Sure there are references in there to things, but LLMs don’t capture that. A Generative Pre-trained Transformer – or any mechanism for generative prediction – is not “understanding” that a cat has four legs, sharp teeth and fur, what it is learning is that we pat cats, we attend when they catch mice, and that they jump on window sills and knock things off. The things we notice about cats are all integrated into our own practices, and what we talk about is those practices.
The hypothesis is that an LLM, rather than abstracting from “cereal” to “food” and then reasoning about “hunger,” is generalising from N descriptions of a practice to a rewriting of the practice selected and modified by the current prompt. Prompted by “I’m hungry” an LLM will “select” the practice of mother feeding a child - because this is a standard opening for a child to a mother - and fill it in. If the prompt includes the statement that the child likes cereal, then “want some cereal” would be a very plausible magic response.
@@ 6 Summary
In large part the achievements of us humans are a product, not of individuals thinking about stuff, but of societies evolving to do what they do. A society is a set of practices that are mutually dependent and supportive; humans are the mechanism by which practices are executed. Mostly. Societies compete, and those with the better combination of practices survive. Of course modern humans can think about and change practices, and our super power is to do that with symbols, but the symbols – critically the signified referenced by the sign – needs to be identified first, and that process is grounded in the practices a society has evolved. Some practices have obvious advantage when looking at the survival of a society. Farming, child rearing and building shelters are fine, but others – going to church or a football match, elections, and street parties – are less clear and indeed it is hard to think about the consequences of doing them differently.
At all levels an individual acts a role in different practices at different times, and does not need to be aware or think about why the practice is good for anyone or anything. What the individual does need to do is to learn how to perform that role and we do that by copying. Whatever that mechanism is, it enables us to predict what happens next. This is how language works. We see a child starting the practice of getting something to eat and we all expect – including the child – that the result will be some food. When that doesn’t happen, we could reason about the contents of the mind of the mother, but we might also have learnt the practice of negotiation.

### v6

@@ Abstract
This paper proposes a model of what Large Language Models are doing when they impress us with its language skills. Rather than doing classic AI with “distributed representations,” the proposal is that LLMs work because they have been trained on descriptions of practices, and “glorified auto-complete” is sufficient to capture the mechanism behind a significant proportion of our rather shallow intelligence. Although not flattering to our modern ego, the approach provides an explanation of apparent “mind reading” in human communication – a phenomena that has been challenging for conversational user interfaces.
@@
For the last few years work in AI has become synonymous with machine learning and in particular neural net architectures for Large Language Models or LLMs. These systems do impressive things, but how they do them is poorly understood. The idea that we do things by thinking about it is very appealing and, with the growth of interest in AI, many coming from outside the field have not studied the history. Inevitably old mistakes get repeated. The paper starts with a recap of good old fashioned AI (GOFAI). It then describes what Brooks and many philosophers have said about the nature of situated action. It then introduces a mechanism that was touched on in the 1970s in AI research but which is an idea that is hundreds of years old in what is now sociology. Finally the paper discusses how the mechanism can be applied to a major issue in Natural Language Understanding (NLU) namely the phenomenon of apparent “mind reading”.
@@ 1 The problem: people read minds
Working on NLU, one sooner or later comes across a phenomenon in which humans recognise the intent of their conversational partner (for example Wallis (2008)). The solution is often assumed to be a better (and more detailed) representation of the context and from that perspective this paper proposes representing context, not as things and their relationships, but as an ecology of practices.
@@ 16 Mother: You want some cereal?
All English speakers by definition will follow what is happening in this conversation and feel there is nothing extraordinary about it. The problem for a reference model of language understanding is that the mother says nothing related to the child’s utterances until line 11. There is no overlap of semantic content - none at all. Of course, as English speakers, we can all say what both the mother and the child want.
This has historically lead researchers interested in conversational agents off into the realm of recognition of intent and modelling theory-of-mind (ToM) Allen et al. (1995); Levinson (2006). Classically this would be modelled in terms of the beliefs, desire and intentions of an agent, the theory being that a rational agent will do what it believes is in its interests Cohen and Levesque (1990), all in terms of good old fashioned symbolic representations. But is that how minds actually work?
@@ 2 Artificial Intelligence - a history
Computers are very good - indeed ideal - at manipulating symbols. In general, words and symbols are seen as having meaning by referring to things in the world, and the early AI researchers saw language understanding as a problem of mapping messy natural languages into something more tidy Mel’cuk (1981) – more pure – as God would have made it Eco (1995). For centuries there has been the idea that meaning bottoms out at some kind of semantic primitives Ishiguro (1990) and in this tradition Roger Schank proposed that there was a workable set of primitive verbs or ‘acts’ on which the meaning of a sentence could be based Schank (1972). Examples are PTRANS for “physical transfer” and MTRANS for mental or knowledge transfer, and that machines could translate natural descriptions into (representations of) meaning and figure out consequences from there.
The primitives approach to AI didn’t really work but there was also the idea that a formal system of symbol manipulations might map isomorphically onto events in the world. In “Godel, Escher, Bach”, Hofstader Hofstader (1979) gives a beautiful example with his pq- system which is a set of syntactic transforms on strings of ’p’, ’q’, and ’-’. The system produces an infinite set of strings including “-p--q---” and “--p--q----” but not “-p---q--”. The number of ’-’s in produced strings just happens to map onto the number of apples in a box as you add more, or take them out. The symbol ’p’ seems to mean “plus” and ’q’, “equals” because they appear in the system of productions in an isomorphic relationship with our system of adding and subtracting. They have no meaning in themselves without the system in which they participate. Rather than meaning being built up by combining primitive features with innate meaning, the “primitives” get meaning from the structure in which they occur.
@@
Computers are good at this kind of syntactic symbol manipulation, and between that and the observation that it is hard to think of a thousand things you know a hundred things about, this kind of reasoning formed the basis of thinking about computers and meaning. Today a spreadsheet is a great way to automate reasoning about some things but it turns out there is a wide range of tasks that fail when handled this way. Counting apples in a box might seem easy, but a box might have 15 apples in it according to Waitrose, 25 according to Aldi, and a horse would eat the lot no matter how many are rotten. Modelling the world with referential symbols is hard because the referent can change. If the idea is to map a room before you vacuum it, cats and people tend move and re-set the model. In the early days of robot football the machines would wait for the ball to stop before deciding what to do. Air traffic control has a similar problem. At the time of writing aircraft have been grounded while air traffic controllers sort out why an aircraft, represented by a flight number, occurred in two locations. There is something wrong with the system’s representation of the state of the world. There were projects set up to formally encode all the data in the world, the CYC project Lenat et al. (1986) being the most explicit. The feeling in the late 1980s was that no matter how powerful computers might become in the future, simply getting the data for symbolic representations was going to be problematic, but perhaps there was a better way.
@@
How this pans out for machines is still not clear and Brooks’ robots are often dismissed as being just “insect level intelligence” with robot developers continuing to add symbolic representations in order to support higher level intelligence (for example see Arkin (1998)). Philosophers on the other hand continue to push the limits of what higher level behaviours can be explained in terms of reactive systems Hendriks-Jansen (1996); Ward et al. (2017); Gallagher (2020); Bar-On (2021). The notion of a “reactive system” describes a mechanism used for AI problems, but it is not an interesting “algorithms and data-structures” solution. The ELIZA mechanism takes a set of if-then rules to map input to output, and applied to the DOCTOR script was found to be quite engaging. That same mechanism is used today for a myriad of chatbots but at the time and for many years after, AI researchers would consider it just a trick.
@@ 3 Natural Intelligence - a theory
We humans are not entirely reactive agents because we can, on occasion, think about things like whether I can afford that whiskey using an expression such as “£38 - £44 = …” but we also have another trick. We humans modify the environment to fit with us. In the cooking gnocchi for lunch scenario, none of the things I interact with are naturally occurring – someone made them. What is more, these things were put where I would find them. I might use insect level intelligence to eat lunch, but surely others must have had thoughts (with symbols) in order to create our benign environment. It seems this is not the case.
Recently we have been talking with sociologists who look at human decision making at the macro level. The strong reductionist position is alive and well in sociology and many take the line that collective action is simply made up from the behaviour of individuals making their own decisions - basically that sociology is just psychology on a massive scale in the same way as drunk physicists will claim chemistry is just physics. Social simulation based on the behaviour of individuals is of course entirely possible with today’s computers. There are however problems with this in that collective action does not fit with rational self interest of individuals. Why did Britain leave the EU? And why does Germany want to go to war with a nation that 2 years ago was a key trading partner? At the other end of the “structure and agency” debate wikipedia is the idea that society consists essentially of structures such as institutions, traditions, and norms, and that people are merely acting out their role in those structures. Post The Enlightenment we collectively like to think we can reason about things and make rational decisions that are better than the status quo. And it is of course not very flattering to be told you are merely a cog in a machine, but a key finding of AI research over the last 70 years is of course that rationality is not all it is cracked up to be.
@@
And like an ecology in the natural environment, there is a certain amount of autopoiesis resulting in an ecology of practices being robust. In the animal kingdom, if one species is suddenly removed a gap appears in the resource/consumer space. The gap is either quickly filled by the remnants of that species, something else comes along to fill the space – rabbits replace bilbies – or the resources are appropriated by neighbours in that space. Reduce the number of cats in Sydney, and brown snakes move in to eat the rats and mice ABC (Australia). If the local supermarket closes, I do not starve, but engage another no doubt older hunter/gatherer practice. I wander about attentively until I find indicators of something to eat - an Aldi sign perhaps - which triggers the more modern practice of shopping.
@@
4 What LLMs are actually doing
Humans go through life with a set of practices that they know how to do. Rather than sensing the world, forming a representation of the salient parts, and then doing (symbolic) reasoning to produce action that happens to string together to form practices, the basic unit of intelligence as we know it is the practice. Practices as part of an ecology of practices produced by a culture and which facilitates the propagation of the culture.
The proposal is that the things we write down are to a very large extent descriptions of practices. LLM training data is a description of human practice. Sure there are references in there to things, but LLMs don’t capture that. A Generative Pre-trained Transformer – or any mechanism for generative prediction – is not “understanding” that a cat has four legs, sharp teeth and fur, what it is learning is that we pat cats, we attend when they catch mice, and that they jump on window sills and knock things off. The things we notice about cats are all integrated into our own practices, and what we talk about is those practices. In order to play chess Dynomight ; Dynomight (updated) LLMs do not represent pieces and strategies; what they do is use “glorified auto complete” Rothman (2023) on descriptions of chess games – on descriptions of the practice of playing chess. It certainly looks like it’s reasoning, but that is just anthropomorphism about how we think we solve the problem.
@@ 5 A ’pragmatics first’ model of NLU
We animals behave in accordance with both the reactive framework in which the environment “causes” action, and a framework of practices which evolved through evolutionary pressures on our society. So how does this relate to the mechanism of language as interaction? Language can of course be used to describe practices in terms of things as described above, but using language is also part of practices.
Returning to the problem of mind reading, the radical enactivists have claimed that in many cases what looks like reasoning about other minds and ToM is actually the “direct perception” of another’s intent. (see discussion of Hutto in Gallagher (2020)) The hypothesis is that an LLM, rather than abstracting from “cereal” to “food” and then reasoning about “hunger,” is generalising from N descriptions of a practice to a rewriting of the practice selected and modified by the current prompt. Prompted by “I’m hungry” an LLM will “select” the practice of mother feeding a child - because this is a standard opening for a child to a mother - and fill it in. If the prompt includes the statement that the child likes cereal, then “want some cereal” would be a very plausible magic response.
For those who know their history of AI, practices may look much like Schank and Abelson’s scripts Schank (1989). The problem was that scripts as envisaged needed to be hand crafted and relied on symbolic representations of things like restaurants, bills, eating and payment. The proposal is that LLMs get this data from internet scrapers. Rather than forming representations, they are using “auto complete” on the raw data. The child in our example dialog above can want to eat something and, from past experience know that making the sounds for "I’m hungry" often results in mother feeding him. We might think that the child means the words he says, and he might, but the act of saying it causes the world to change such that there is no longer a sensation of hunger.
@@
6 Human scale intelligence
How does a head learn practices? The proposal is that a head does good old fashioned unsupervised learning on the raw data of what actions lead to which opportunities. The world, in combination with the human body’s ability to act on it, provides interaction that is consistent enough for regularities to be detected. As a consequence, the mind has a catalog of "sequences of interaction" - of practices - which are coherent enough to provide something with a beginning, a middle, and an end. When the agent/world interaction triggers a practice beginning, the agent can predict that, should the middle follow, the end will occur. If the agent can provide that middle, and somehow wants the outcome, then the agent can do what looks like rational action without using symbolic representations of anything.
What the individual does need to do is to learn how to perform roles but that can done without representation. The individual no doubt can learn by directly experiencing sequences of action and reaction, but many animals, including humans, also seem to experience the experiences of others (“second person inter-subjectivity” – see Gallagher (2020)). We also imitate the practices of others (see for example Gergely and Csibra Gergely and Csibra (2006)). There are certainly interesting species specific behaviours that enable the training of individuals to perform roles but the key is that these often operate at a very shallow level.
Key to the mechanism as communication is that practices are largely shared by the community of speakers, and that the practices are largely visible to others. If someone walks up to an ATM machine with a card in their hand, it is fairly likely that they will not start ordering a coffee. This is obvious to us humans – we ‘directly perceive,’ not just the actions, but also the trajectory of which the actions are the start. And it is not just humans. A sheep dog does not need to model the mind of a sheep – the sheep’s observable actions are simply a recognisable part of a practice that sheep perform (Figure 2).
Figure 2: Zak doesn’t read the Ewe’s mind, but instead sees a nasty head but unfolding.
@@ 7 Summary
In large part the achievements of us humans are a product, not of individuals thinking about stuff, but of societies evolving to do what they do. A society is a set of practices that are mutually dependent and supportive; human agents are merely the mechanism by which practices are executed. Mostly. Societies compete, and those with the better combination of practices survive. Of course modern humans can think about and change practices, and our super power is to do that with symbols, but the symbols – critically the signified referenced by the sign – needs to be identified first. That process is grounded in the practices a society has evolved. Some practices have obvious advantage when looking at the survival of a society. Farming, child rearing and building shelters are fine, but others – going to church or a football match, elections, and street parties – are less clear.
At all levels an individual acts a role in different practices at different times, and does not need to be aware or think about why the practice is good for anyone or anything. Whatever that mechanism, representing practices enables us to predict what happens next. This is how language works. We see a child starting the practice of getting something to eat and we all expect – including the child – that the result will be some food. When that doesn’t happen, we could reason about the contents of the mind of the mother but we might, like the dog above, simply recognise the consequences of a power relationship.